# library.bib

### v1

@@ 3.3 Gasket decomposition
As derived by Borot in [borot2014formal], planar stuffed maps have a nested structure, which can be seen in the form of a functional relation which we will now briefly review. Recall the gasket $M^{\prime}$ of a rooted stuffed map $M$ is the rooted ordinary map obtained from $M$ by removing all branches and only keeping the connected component with the root. The resulting map will have some faces that come from 2-cells with one boundary and some that come from multi-boundary 2-cells. In the latter case, we call the faces cement faces. Since $M$ is planar, the cement faces must all be from distinct 2-cells, otherwise such a 2-cell would create a handle in $M$. The original stuffed map $M$ can be reconstructed from the gasket $M^{\prime}$ by replacing all the cement faces with their original 2-cells. At the level of generating functions, this corresponds to replacing the weights of the cement faces with a weighted combination of $\mathcal{T}_{\ell}$.
@@ 3.4 Tutte’s recursion for stuffed maps
This recursive formula can be obtained by applying the functional relation (7) (8) to Tutte’s recursion for ordinary maps (5). See the beginning of Section 4.1 of [borot2014formal] for a graphical derivation.
@@
As noted in [borot2014formal], the functional equation combined with the factorization of the discriminant in the case of ordinary maps implies the same type of factorization for stuffed maps. Once again, the Zhukovsky method can be used to solve for an explicit solution. However, the problem is more complicated to solve in practice due to the appearance of the generating functions $\mathcal{T}_{\ell}^{\prime}s$ in $\mathcal{V}(x)$. For an example of such a solution, see [hessam2023double].
@@ 3.5 The tree equation for stuffed maps with pointed gaskets
We will now apply the tree equation for ordinary maps and the functional equation 3 to show that $\gamma^{2}$ for stuffed maps is the root of an equation.
@@
Note that one may write each generating function for ordinary maps strictly in terms of $\gamma$ via Corollary 3.1.2 of [eynard2016counting] which states
@@
If the set of stuffed maps is glued from a finite set of 2-cells, then this equation is a polynomial and $\gamma$ is an algebraic. In its full generality, it is not entirely clear to the author how to bijectively interpret this equation. However, we will discuss this bijection in detail for a specific case in Section 4. Note that by using the above proposition, Corollary 3.1.2 of [eynard2016counting], and Borot’s functional equation, one can explicitly compute any generating function $\mathcal{T}_{\ell}$ in terms of the Boltzmann weights. This will be done for an example in Section 4.
@@ 3.6 Adding a point
Proposition 3.
The following functional relationship holds between the generating function of pointed stuffed maps with boundaries and the generating function of rooted stuffed maps with pointed gaskets:
@@
Recall Lemma 3.1.4 from [eynard2016counting], which gives us simplified forms of the derivatives of the generating function $W(x)$ in Zhukovsky space
@@ 4 Example: Bridged quadrangulations
Bridged quadrangulations appear in the computation of a quartic matrix model with a $(\operatorname{Tr}H^{2})^{2}$ interaction [das1990new, klebanov1995non, hessam2022noncommutative, khalkhali2024coloured]. Let $S$ consist of a quadrangle and a 2-cell with two boundaries of length 2. We refer to the latter 2-cell as a bridge and the maps glued from $S$ (as well as some boundaries of fixed lengths) as bridged quadrangulations.
@@
[…]

### v2

@@ 3.3 Gasket decomposition
As derived by Borot in [5], planar stuffed maps have a nested structure, which can be seen in the form of a functional relation which we will now briefly review. Recall the gasket $M^{\prime}$ of a rooted stuffed map $M$ is the rooted ordinary map obtained from $M$ by removing all branches and only keeping the connected component with the root. The resulting map will have some faces that come from 2-cells with one boundary and some that come from multi-boundary 2-cells. In the latter case, we call the faces cement faces. Since $M$ is planar, the cement faces must all be from distinct 2-cells, otherwise such a 2-cell would create a handle in $M$. The original stuffed map $M$ can be reconstructed from the gasket $M^{\prime}$ by replacing all the cement faces with their original 2-cells. At the level of generating functions, this corresponds to replacing the weights of the cement faces with a weighted combination of $\mathcal{T}_{\ell}$.
@@ 3.4 Tutte’s recursion for stuffed maps
This recursive formula can be obtained by applying the functional relation (7) (8) to Tutte’s recursion for ordinary maps (5). See the beginning of Section 4.1 of [5] for a graphical derivation.
@@
As noted in [5], the functional equation combined with the factorization of the discriminant in the case of ordinary maps implies the same type of factorization for stuffed maps. Once again, the Zhukovsky method can be used to solve for an explicit solution. However, the problem is more complicated to solve in practice due to the appearance of the generating functions $\mathcal{T}_{\ell}^{\prime}s$ in $\mathcal{V}(x)$. For an example of such a solution, see [17].
@@ 3.5 The tree equation for stuffed maps with pointed gaskets
We will now apply the tree equation for ordinary maps and the functional equation 3.1 to show that $\gamma^{2}$ for stuffed maps is the root of an equation.
@@
Note that one may write each generating function for ordinary maps strictly in terms of $\gamma$ via Corollary 3.1.2 of [14] which states
@@
If the set of stuffed maps is glued from a finite set of 2-cells, then this equation is a polynomial and $\gamma$ is an algebraic. In its full generality, it is not entirely clear to the author how to bijectively interpret this equation. However, we will discuss this bijection in detail for a specific case in Section 4. Note that by using the above proposition, Corollary 3.1.2 of [14], and Borot’s functional equation, one can explicitly compute any generating function $\mathcal{T}_{\ell}$ in terms of the Boltzmann weights. This will be done for an example in Section 4.
@@ 3.6 Adding a point
Theorem 3.1.
The following functional relationship holds between the generating function of pointed stuffed maps with boundaries and the generating function of stuffed maps with pointed gaskets:
@@
Recall Lemma 3.1.4 from [14], which gives us simplified forms of the derivatives of the generating function $W(x)$ in Zhukovsky space
@@ 4 Example: Bridged quadrangulations
Bridged quadrangulations appear in the computation of a quartic matrix model with a $(\operatorname{Tr}H^{2})^{2}$ interaction [12, 19, 16, 18]. Let $S$ consist of a quadrangle and a 2-cell with two boundaries of length 2. We refer to the latter 2-cell as a bridge and the maps glued from $S$ (as well as some boundaries of fixed lengths) as bridged quadrangulations.
@@
$\gamma^{2}=\frac{1}{2}\sqrt{Q-\frac{R}{3\sqrt[3]{2}t_{2,2}t_{4}}}-\frac{1}{2}\sqrt{\frac{4}{t_{2,2}t_{4}\sqrt{Q-\frac{R}{3\sqrt[3]{2}t_{2,2}t_{4}}}}+Q+\frac{R}{3\sqrt[3]{2}t_{2,2}t_{4}}},$